OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2x - 1 >= -7? x >= -3 x<= -3 -3 <= x A number line with a closed circle on negative 3 and shading to the right. @ganeshie8

ganeshie8 (ganeshie8):

first strike off the obvious ones

ganeshie8 (ganeshie8):

wat did u get after solving the inequality ?

OpenStudy (jhannybean):

\[\large 2x-1 \ge -7\]\[\large 2x +1-1 \ge -7 +1\]\[\large 2x \ge -6\]\[\large x \ge ?\]

OpenStudy (anonymous):

i got x<=-3

ganeshie8 (ganeshie8):

look at jhanny work... you must get x >= -3

OpenStudy (jhannybean):

What @ganeshie8 meant was to cancel out the options that had the sign flipped. You only flip the inequality sign when you're dividing or multiply by the coefficient of the variable you are evaluating for.

OpenStudy (anonymous):

so then the answer would be B?

ganeshie8 (ganeshie8):

brilliant !

OpenStudy (jhannybean):

good job :)

ganeshie8 (ganeshie8):

oops ! i think you meant to say answer is x <= -3, previously is it

OpenStudy (jhannybean):